# Non-autoregressive Generative Models for Reranking Recommendation

### v1

@@ Abstract.
Contemporary recommendation systems are designed to address the daily requirements of users by delivering a curated list of items that align with the user’s specific demands or interests. In a multi-stage recommendation system, reranking plays a crucial role by modeling the intra-list correlations among items.
The key challenge of reranking lies in the exploration of optimal sequences within the combinatorial space of permutations. Recent research proposes a generator-evaluator learning paradigm, where the generator generates multiple feasible sequences and the evaluator picks out the best sequence based on the estimated listwise score. Generator is of vital importance, and generative models are well-suited for the generator function. Current generative models employ an autoregressive strategy for sequence generation. However, deploying autoregressive models in real-time industrial systems is challenging. Firstly, the generator can only generate the target items one by one and hence suffers from slow inference. Secondly, the discrepancy between training and inference brings an error accumulation. Lastly, the left-to-right generation neglects information from succeeding items, resulting in suboptimal performance.
Hence, we propose a Non-AutoRegressive generative model for reranking Recommendation (NAR4Rec) designed to enhance efficiency and effectiveness. To address challenges related to sparse training samples and dynamic candidates impacting model convergence, we introduce a matching model. Considering the diverse nature of user feedback, we propose a sequence-level unlikelihood training objective to distinguish feasible from unfeasible sequences. Additionally, to overcome the lack of dependency modeling in non-autoregressive models regarding target items, we introduce contrastive decoding to capture correlations among these items.
Extensive offline experiments on publicly available datasets validate the superior performance of our proposed approach compared to the existing state-of-the-art reranking methods. Furthermore, our method has been fully deployed in a popular video app Kuaishou with over 300 million daily active users, significantly enhancing online recommendation quality, and demonstrating the effectiveness and efficiency of our approach.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systemsccs: Computing methods  Generative models
@@ 1. Introduction
Recommendation systems provide users with a personalized list of items that are tailored to their needs by aligning the recommendation with the user’s interest. With the emergence of deep learning, various approaches have been proposed to discern and capture user interests through the analysis of their actions. These models primarily focus on intricate feature interaction(Cheng et al., 2016; Guo et al., 2017; Lian et al., 2018), user preference modeling(Zhou et al., 2018, 2019b) and so on. However, most existing methods focus on modeling individual items separately using an itemwise learning paradigm. However, those methods fail to capture the mutual influence between items, resulting in suboptimal performance. Acknowledging that whether the user interactions with one item may correlate with others in the recommendation list(Pei et al., 2019), the reranking stage is introduced to consider the context information and then generate an optimal sequence of items of recommended items.
@@
The key challenge of reranking lies in the exploration of optimal sequences within the combinatorial space of permutations. Existing methods for reranking can be categorized into one-stage and two-stage methods. One-stage methods(Ai et al., 2018; Pei et al., 2019) take candidates as input, estimating refined scores for each item under the permutation, and then rerank the candidates based on these refined scores in a greedy manner. However, one-stage methods encounter an inherent contradiction(Feng et al., 2021; Xi et al., 2021): the reranking operation inherently alters the permutation, introducing distinct mutual influences between items compared to the initial permutation. Consequently, the refined score conditioned on the initial permutation is considered implausible, indicating a coherence limitation in these one-stage methods.
@@
To overcome this challenge, two-stage methods employ a generator-evaluator framework where the generator produces multiple feasible sequences, and the evaluator selects the optimal sequence based on the estimated listwise score. In the generator-evaluator framework, the generator is of vital importance. Generative models(Bello et al., 2018; Feng et al., 2021; Gong et al., 2022; Zhuang et al., 2018) are preferable over heuristic methods(Feng et al., 2021; Lin et al., 2023; Xi et al., 2021; Shi et al., 2023) for the generator function considering the expansive solution space of item permutations. Generative models utilized in this framework commonly employ an autoregressive strategy for sequence generation. However, deploying the autoregressive models in real-time industrial recommendation systems poses challenges.
Firstly, autoregressive models suffer from inference efficiency. Autoregressive models adopt a sequential approach to generate target sequences item by item, resulting in slow inference. The time complexity scales linearly with the sequence length during generation, leading to high latency in real-time applications.
@@
Secondly, a critical issue arises from the training-inference discrepancy in autoregressive models. While these models are trained to predict the next item based on the ground truth up to that timestep, during inference, they receive their own previously generated outputs as input. This misalignment introduces an accumulated error, where inaccuracies generated in earlier timesteps propagate and accumulate over time. Consequently, this accumulation leads to sequences that deviate from the true distribution of the target sequence.
Additionally, autoregressive models exhibit limited information consideration. The sequential decoding process focuses solely on preceding items, neglecting information from succeeding items. This limitation results in suboptimal performance as the model fails to fully leverage the available context.
@@
Figure 1. Comparison of autoregressive model (left) and non-autoregressive model (right). Auto-regressive models generate items sequentially while non-autoregressive generate all items simultaneously.
@@
To address those challenges, we propose Non-AutoRegressive generative model for reranking Recommendation, referred to as NAR4Rec. As illustrated in fig. 1, traditional autoregressive models adopt a step-by-step approach to generate target sequences, where the model takes its own previously generated outputs as input. Conversely, our proposed NAR4Rec model generates all items within the target sequence simultaneously in a single model forward pass.
However, we find it nontrivial to deploy non-autoregressive in recommendation systems.
  • Firstly, the sparse nature and dynamic candidates in recommendation systems pose difficulties for learning convergence, which we address by sharing position embedding and introducing a matching model.
  • Secondly, the diverse nature of user feedback, including both positive and negative interactions, renders maximum likelihood training less suitable. We propose unlikelihood training to distinguish between desirable and undesirable sequences.
  • Lastly, canonical non-autoregressive models assume independent selection of items at each position in a sequence, which is inadequate when modeling intra-list correlation. We propose contrastive decoding to capture dependencies across items and address this limitation.
@@
  • To the best of our knowledge, we make the first attempt to adopt non-autoregressive models for reranking, which speeds up the inference speed and meets the requirements of real-time recommendation systems.
  • We propose a matching model for better convergence, a sequence-level unlikelihood training method for guiding the generated sequence towards better overall utility, and a contrastive search to constrain current decoding strategies with intra-list correlation.
  • Extensive offline experiments conducted on public dataset demonstrate that NAR4Rec outperforms state-of-the-art methods. Online experiments also shows the effectiveness of NAR4Rec. Furthermore, NAR4Rec have fully deployed in a real-world video app with over 300 million daily active users, significantly improve the recommendation quality.
@@ 2.1. Reranking in Recommendation Systems
One-stage methods treat reranking as a retrieval task, recommending the top k items based on scores from a ranking model. These methods refine the initial list distribution using list-wise information, optimizing overall recommendation quality. Subsequently, the candidates are reranked by the refined itemwise score in a greedy manner. The distinction lies in the network architectures for capturing list-wise information, such as GRU in DLCM(Ai et al., 2018), and transformer in PRM(Pei et al., 2019). However, user feedback for the exposed list is influenced not just by item interest but also by arrangements and surrounding context(Joachims et al., 2017; Lorigo et al., 2008, 2006; Yang, 2017). The reranking operation modifies permutations, thereby introducing influences distinct from the initial permutation. Moreover, ne-stage methods, which exclusively model the initial permutation, fall short in capturing alternative permutations. Consequently, those methods struggle to maximize overall user feedback(Xi et al., 2021).
Two-stage methods (Feng et al., 2021; Shi et al., 2023; Lin et al., 2023; Xi et al., 2021) embrace a generator-evaluator framework. In this framework, the generator initiates the process by generating multiple feasible sequences, and subsequently, the evaluator selects the optimal sequence based on the estimated listwise score. This framework allows for a comprehensive exploration of various feasible sequences, and an informed selection of the most optimal one based on listwise considerations. The role of the generator is particularly crucial for generating sequences. Common approaches for generator can be categorized into heuristic methods(Feng et al., 2021; Lin et al., 2023; Xi et al., 2021; Shi et al., 2023), such as beam search or item swapping, and generative models (Bello et al., 2018; Feng et al., 2021; Gong et al., 2022; Zhuang et al., 2018). Generative models are more suitable than heuristic methods for reranking given the vast permutation space. These generative models typically adopt a step-greedy strategy which autoregressively decides the display results of each position. However, the high computational complexity of online inferences limits their application in real-time recommendation systems.
To address the challenges associated with autoregressive generation models, our work explores the potential of non-autoregressive generative models as the generator in the generator-evaluator framework. Non-autoregressive generative models generate the target sequence once to alleviate computational complexity.
@@ 2.2. Non-autoregressive Sequence Generation
Non-autoregressive sequence generation(Gu et al., 2018) is initially introduced in machine translation to speed up the decoding process. Then non-autoregressive has gained increased attention in nature language processing, e.g. text summarization(Gu et al., 2019; Awasthi et al., 2019), text error correction(Leng et al., 2021b, a). In particular, research efforts have been directed toward addressing challenges arising from the lack of target information in non-autoregressive models. Strategies include enhancing the training corpus to mitigate target-side dependencies(Gu et al., 2018; Zhou et al., 2019a) and refining training approaches(Ghazvininejad et al., 2019; Stern et al., 2019) to alleviate learning difficulties.
Though non-autoregressive generation has been explored in Natural Language Processing, the conventional techniques are not directly applicable to recommendation systems. We tackle the challenges encountered by recommendation to improve the convergence and performance of non-autoregressive models and make the first attempt to integrate non-autoregressive models into reranking in real-time recommender systems.
@@ 3.1. Reranking problem Formulation
For each user $u$ within the set $U$, a request encompasses a set of user profile features(such as user ID, gender, age), the recent interaction history and $n$ candidates items denoted as $X=\left\{x_{1},x_{2},\cdots,x_{n}\right\}$, where $n$ is the number of candidates. Given candidates $X$, the goal of reranking is to propose an item sequence $Y=\left\{y_{1},y_{2},\cdots,y_{m}\right\}$ with the most favorable feedback for user $u$, where $m$ is the sequences length. We denote the reranking models as $\mathcal{F}(u,\theta,X)$ where the corresponding parameter is $\theta$ and $Y$ is the outcome of the reranking model. In real-time recommendation systems, reranking acts as the last stage to deliver the ultimate list of recommended items. Typically, $n$ significantly exceeds $m$, with $m$ being less than 10 and $n$ ranging from several tens to hundreds.
In a multi-interaction scenario, users may exhibit distinct types of interaction (e.g., clicks, likes, comments) for each item exposure. Formally, we define the set of user interactions as $B$, and $y_{u,i,b}$ represents user $u$’s response to item $y_{i}$ concerning interaction $b\in B$. Given ${Y}$, each item $y_{i}$ elicits a multi-interaction response $\mathbf{s}_{u,y_{i}}=[{s}_{u,y_{i},b_{1}},\ldots,{s}_{u,y_{i},b_{|{B}|}}]$ and the overall user response is:
(1) $\mathbf{S}_{u,Y}=\left[\begin{matrix}s_{u,y_{1},b_{1}}&\dots&s_{u,y_{m},b_{1}}\\ \vdots&\ddots&\vdots\\ s_{u,y_{1},b_{|B|}}&\dots&s_{u,y_{m},b_{|B|}}\end{matrix}\right].$
The overall utility is quantified as the summation of individual item utilities, denoted as $\mathcal{R}(u,Y)=\sum_{i=1}^{m}\mathcal{R}(u,y_{i})$. The utility associated with each item may correspond to a specific interaction type $b$, such as clicks, watch time, or likes. In such cases, the item utility is expressed as $\mathcal{R}(u,y_{i})=s_{u,y_{i},b}$. Alternatively, the item utility can be represented as the weighted sum of diverse interactions $\mathcal{R}(u,y_{i})=\sum_{b}w_{b}s_{u,y_{i},b}$, where $w_{b}$ denotes the weight for each interaction.
@@
The reranking phase introduces a permutation space with exponential size, precisely expressed as $\mathcal{O}(A_{n}^{m})$. This denotes the solution space for selecting and ordering $m$ items from a pool of $n$ candidates. Considering that items in a list have intricate connections to each other, users display unique feedback for each permutation. However, in practical scenarios, users are exposed to only one permutation. Consequently, the primary challenge in reranking lies in efficiently and effectively determining the permutation-optimal list given the vast solution space yet extremely sparse real user feedback as training samples.
@@ 3.2. Autoregressive sequence generation
(3) $p_{\mathcal{\text{AR}}}(Y|X;\theta)=\sum_{t=1}^{m+1}p(y_{t}|y_{0:t-1},x_{1:n};\theta).$
where the special tokens $y_{0}$ (e.g., ¡bos¿) and $y_{m+1}$ (e.g., ¡eos¿) denote the beginning and end of target sequences. Importantly, the length of the generated sequence is predetermined and fixed, unlike variable lengths in natural language generation.
@@
(4) $\mathcal{L}_{\text{AR}}=-\log p_{\text{AR}}(Y|X;\theta)=-\sum_{t=1}^{m+1}\log p(y_{t}|y_{0:t-1},x_{1:n};\theta).$
The training objective aims to optimize individual conditional probabilities. During training, where the target sequence is known, the computation of these probabilities is conditioned on the target items at earlier timesteps rather than the model-generated output items, allowing for efficient parallelization. During inference, the autoregressive models generate the target sequence item-by-item sequentially. Autoregressive models align well with the inherent browsing nature of recommendation systems, efficiently capturing the distribution of the target sequence. This makes them well-suited for the reranking task, particularly considering the vast space of possible permutations.
While autoregressive models have demonstrated effectiveness, their deployment in industrial recommendation systems is challenging. Firstly, their sequential decoding nature leads to slow inference, introducing latency that hinders real-time application. Secondly, these models, trained to predict based on ground truth, face a discrepancy during inference, receiving their own generated outputs as input. This misalignment may lead to compounded errors, as inaccuracies generated in earlier timesteps accumulate over time, resulting in inconsistent or divergent sequences that deviate from the true distribution of the target sequence. Thirdly, autoregressive models rely on a left-to-right causal attention mechanism, limiting the expressive power of hidden representations, as each item encodes information solely from preceding items(Sun et al., 2019). This constraint impedes optimal representation learning, resulting in suboptimal performance.

### v3

@@ Abstract.
Contemporary recommendation systems are designed to meet users’ needs by delivering tailored lists of items that align with their specific demands or interests. In a multi-stage recommendation system, reranking plays a crucial role by modeling the intra-list correlations among items. The key challenge of reranking lies in the exploration of optimal sequences within the combinatorial space of permutations. Recent research proposes a generator-evaluator learning paradigm, where the generator generates multiple feasible sequences and the evaluator picks out the best sequence based on the estimated listwise score. The generator is of vital importance, and generative models are well-suited for the generator function. Current generative models employ an autoregressive strategy for sequence generation. However, deploying autoregressive models in real-time industrial systems is challenging. Firstly, the generator can only generate the target items one by one and hence suffers from slow inference. Secondly, the discrepancy between training and inference brings an error accumulation. Lastly, the left-to-right generation overlooks information from succeeding items, leading to suboptimal performance.
To address these issues, we propose a Non-AutoRegressive generative model for reranking Recommendation (NAR4Rec) designed to enhance efficiency and effectiveness. To tackle challenges such as sparse training samples and dynamic candidates, we introduce a matching model. Considering the diverse nature of user feedback, we employ a sequence-level unlikelihood training objective to differentiate feasible sequences from unfeasible ones. Additionally, to overcome the lack of dependency modeling in non-autoregressive models regarding target items, we introduce contrastive decoding to capture correlations among these items. Extensive offline experiments validate the superior performance of NAR4Rec over state-of-the-art reranking methods. Online A/B tests reveal that NAR4Rec significantly enhances the user experience. Furthermore, NAR4Rec has been fully deployed in a popular video app Kuaishou with over 300 million daily active users.
@@
$\dagger$ Corresponding author
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining ; August 25–29, 2024; Barcelona, Spain.booktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spainisbn: 979-8-4007-0490-1/24/08doi: 10.1145/3637528.3671645ccs: Information systems Retrieval models and rankingccs: Computing methodologies Machine learning algorithms
@@ 1. Introduction
Recommendation systems offer users personalized item lists tailored to their interests. Various approaches have been proposed to capture user interests, focusing on feature interactions(Cheng et al., 2016; Guo et al., 2017; Lian et al., 2018), user preference modeling(Zhou et al., 2018, 2019b), and so on. However, most existing methods treat individual items separately, neglecting their mutual influence and leading to suboptimal results. Acknowledging that user interactions with one item may correlate with others in the recommendation list(Pei et al., 2019), reranking is introduced to consider contextual information and generate an optimal sequence of recommendation items.
@@
The main challenge in reranking is exploring optimal sequences within the vast space of permutations. Reranking methods are typically categorized into on-stage and two-stage approaches. One-stage methods (Ai et al., 2018; Pei et al., 2019) take candidates as input, estimating refined scores for each item within the permutation, and rerank them greedily based on these scores. However, one-stage methods encounter an inherent contradiction(Feng et al., 2021; Xi et al., 2021): the reranking operation inherently alters the permutation, introducing different mutual influences between items compared to the initial arrangement. Consequently, the refined score conditioned on the initial permutation is considered implausible.
@@
To tackle this challenge, two-stage methods utilize a generator-evaluator framework. Here, the generator creates multiple feasible sequences, and the evaluator selects the optimal sequence based on the estimated listwise score. Within the generator-evaluator framework, the generator plays a crucial role. Generative models(Bello et al., 2018; Feng et al., 2021; Gong et al., 2022; Zhuang et al., 2018) are preferred over heuristic methods(Feng et al., 2021; Lin et al., 2023; Xi et al., 2021; Shi et al., 2023) for the generator function due to the expansive solution space of item permutations. Generative models commonly employ an autoregressive strategy for sequence generation.
@@
However, deploying the autoregressive models in real-time industrial recommendation systems presents challenges. Firstly, autoregressive models suffer from inference efficiency. Autoregressive models adopt a sequential approach to generate target sequences item by item, resulting in slow inference as the time complexity increases linearly with the sequence length. Secondly, a critical issue arises from the training-inference discrepancy in autoregressive models. While these models are trained to predict the next item based on the ground truth up to that point. However, during inference, they receive their own previously generated outputs as input. This misalignment introduces an accumulated error, where inaccuracies generated in earlier timesteps propagate and accumulate over time. Consequently, this accumulation leads to sequences that deviate from the true distribution of the target sequence. Additionally, autoregressive models have limited information utilization. The sequential decoding process focuses solely on preceding items, neglecting information from succeeding items. This limitation results in suboptimal performance as the model fails to fully leverage the available context.
@@
Figure 1. Comparison between autoregressive model (left) and non-autoregressive model (right). Auto-regressive models generate items sequentially while non-autoregressive generate all items simultaneously.
@@
To address those challenges, we propose Non-AutoRegressive generative model for reranking Recommendation(NAR4Rec). Unlike autoregressive models, which generate sequences step by step, relying on their own previous outputs, NAR4Rec generates all items within the target sequence simultaneously.
However, we find it nontrivial to deploy non-autoregressive in recommendation systems. Firstly, the sparse nature and dynamic candidates in recommendation systems pose difficulties for learning convergence, which we address by sharing position embedding and introducing a matching model. Secondly, the diverse nature of user feedback, including both positive and negative interactions, renders maximum likelihood training less suitable. We propose unlikelihood training to distinguish between desirable and undesirable sequences. Lastly, non-autoregressive models assume an independent selection of items at each position in a sequence, which is inadequate when modeling intra-list correlation. Hence we propose contrastive decoding to capture dependencies across items.
@@
  • We make the first attempt to adopt non-autoregressive models for reranking, which significantly speeds up the inference speed and meets the requirements of real-time recommendation systems.
  • We propose a matching model to enhance convergence, a sequence-level unlikelihood training method to guide the generated sequence towards improved overall utility, and a contrastive decoding method to refine current decoding strategies with intra-list correlation.
  • Extensive offline experiments demonstrate that NAR4Rec outperforms state-of-the-art methods. Online A/B tests further validate the effectiveness of NAR4Rec. Furthermore, NAR4Rec has been fully deployed in a real-world video app Kuaishou with over 300 million daily active users, notably improving the user experience.
@@ 2.1. Reranking in Recommendation Systems
One-stage methods treat reranking as a retrieval task, recommending the top k items based on scores from a ranking model. These methods refine the initial list distribution using list-wise information, optimizing overall recommendation quality. Subsequently, the candidates are reranked by the refined itemwise score in a greedy manner. The distinction lies in the network architectures for capturing list-wise information, such as GRU in DLCM(Ai et al., 2018), and transformer in PRM(Pei et al., 2019). However, user feedback for the exposed list is influenced not just by item interest but also by arrangements and surrounding context(Joachims et al., 2017; Lorigo et al., 2008, 2006; Yang, 2017). The reranking operation modifies permutations, thereby introducing influences distinct from the initial permutation. Moreover, one-stage methods, which exclusively model the initial permutation, fall short of capturing alternative permutations. Consequently, those methods struggle to maximize overall user feedback(Xi et al., 2021).
Two-stage methods (Feng et al., 2021; Shi et al., 2023; Lin et al., 2023; Xi et al., 2021) embrace a generator-evaluator framework. In this framework, the generator initiates the process by generating multiple feasible sequences, and subsequently, the evaluator selects the optimal sequence based on the estimated listwise score. This framework allows for a comprehensive exploration of various feasible sequences, and an informed selection of the most optimal one based on listwise considerations. The role of the generator is particularly crucial for generating sequences. Common approaches for generators can be categorized into heuristic methods(Feng et al., 2021; Lin et al., 2023; Xi et al., 2021; Shi et al., 2023), such as beam search or item swapping, and generative models (Bello et al., 2018; Feng et al., 2021; Gong et al., 2022; Zhuang et al., 2018). Generative models are more suitable than heuristic methods for reranking given the vast permutation space. These generative models typically adopt a step-greedy strategy which autoregressively decides the display results of each position. However, the high computational complexity of online inferences limits their application in real-time recommendation systems.
To address the challenges linked with autoregressive generation models, our work investigates the viability of non-autoregressive generative models within the generator-evaluator framework. Non-autoregressive generative models generate the target sequence once to alleviate computational complexity.
@@ 2.2. Non-autoregressive Sequence Generation
Non-autoregressive sequence generation(Gu et al., 2018) was initially introduced in machine translation to speed up the decoding process. Then it has since gained increased attention in nature language processing, e.g. text summarization(Gu et al., 2019; Awasthi et al., 2019), text error correction(Leng et al., 2021b, a). Specifically, efforts have been focused on tackling the absence of target information in non-autoregressive models. Strategies include enhancing the training corpus to mitigate target-side dependencies(Gu et al., 2018; Zhou et al., 2019a) and refining training approaches(Ghazvininejad et al., 2019; Stern et al., 2019) to alleviate learning difficulties.
Although non-autoregressive generation has been explored in text, those conventional techniques are not directly applicable to recommendation systems. We tackle the challenges encountered in recommendation systems to improve the convergence and performance of non-autoregressive models and make the first attempt to integrate non-autoregressive models into reranking within real-time recommender systems.
@@ 3.1. Reranking problem Formulation
For each user $u$ within the set $U$, a request encompasses a set of user profile features(such as user ID, gender, age), the recent interaction history, and $n$ candidates items denoted as $X=\left\{x_{1},x_{2},\cdots,x_{n}\right\}$, where $n$ is the number of candidates. Given candidates $X$, the goal of reranking is to propose an item sequence $Y=\left\{y_{1},y_{2},\cdots,y_{m}\right\}$ that elicits the most favorable feedback for user $u$, where $m$ is the sequence length and $Y$ is the recommended list of the reranking model. We denote the reranking models as $\mathcal{F}(u,\theta,X)$ where the corresponding parameter is $\theta$. In real-time recommendation systems, reranking acts as the last stage to deliver the ultimate list of recommended items. Typically, $n$ significantly exceeds $m$, with $m$ being less than 10 and $n$ ranging from several tens to hundreds.
In a multi-interaction scenario, users may exhibit distinct types of interaction (e.g., clicks, likes, comments) for each item exposure. Formally, we define the set of user interactions as $B$, and $s_{u,y_{i},b}$ represents user $u$’s response to item $y_{i}$ concerning interaction $b\in B$. Given ${Y}$, each item $y_{i}$ has a multi-interaction response $\mathbf{e}_{u,y_{i}}=[\textnormal{e}_{u,y_{i},b_{1}},\ldots,\textnormal{e}_{u,y_{i},b_{|{B}|}}]$. For all items ${Y}$,the overall user response is:
(1) $\mathbf{E}_{u,Y}=\left[\begin{matrix}e_{u,y_{1},b_{1}}&\dots&e_{u,y_{m},b_{1}}\\ \vdots&\ddots&\vdots\\ e_{u,y_{1},b_{|B|}}&\dots&e_{u,y_{m},b_{|B|}}\end{matrix}\right].$
The overall utility is quantified as the summation of individual item utilities, denoted as $\mathcal{R}(u,Y)=\sum_{i=1}^{m}\mathcal{R}(u,y_{i})$. The utility associated with each item may correspond to a specific interaction type $b$, such as clicks, watch time, or likes. In such cases, the item utility is expressed as $\mathcal{R}(u,y_{i})=e_{u,y_{i},b}$. Alternatively, the item utility can be represented as the weighted sum of diverse interactions $\mathcal{R}(u,y_{i})=\sum_{b}w_{b}e_{u,y_{i},b}$, where $w_{b}$ denotes the weight for each interaction.
@@
Reranking introduces a permutation space with exponential size, represented as $\mathcal{O}(A_{n}^{m})$, where $n$ represents the number of candidates and $m$ represents the number of items to be selected and ordered. Each permutation represents a potential arrangement of items, and users provide unique feedback for each permutation. However, in practical scenarios, users typically encounter only one permutation. Thus, the main challenge in reranking lies in efficiently and effectively determining the optimal permutation given the vast solution space yet extremely sparse real user feedback as training samples.
@@ 3.2. Autoregressive sequence generation
(3) $p_{\mathcal{\text{AR}}}(Y|X;\theta)=\sum_{i=1}^{m+1}p(y_{i}|y_{0:i-1},x_{1:n};\theta),$
where the special tokens $y_{0}$ (e.g., ¡bos¿) and $y_{m+1}$ (e.g., ¡eos¿) denote the beginning and end of target sequences. Importantly, the length of the generated sequence is predetermined and fixed, unlike variable lengths in text.
@@
(4) $\mathcal{L}_{\text{AR}}=-\log p_{\text{AR}}(Y|X;\theta)=-\sum_{i=1}^{m+1}\log p(y_{i}|y_{0:i-1},x_{1:n};\theta).$
The training objective aims to optimize individual conditional probabilities. In training, when the target sequence is known, these probabilities are calculated based on earlier target items rather than model-generated ones, enabling efficient parallelization. In inference, autoregressive models generate the target sequence item-by-item sequentially, efficiently capturing the distribution of the target sequence. This makes them well-suited for the reranking task, particularly considering the vast space of possible permutations.
While autoregressive models have proven effective, deploying them in industrial recommendation systems is challenging. Firstly, their sequential decoding process leads to slow inference, introducing latency that hinders real-time application. Secondly, these models, trained to predict based on ground truth, face a discrepancy during inference when they receive their own generated outputs as input. This misalignment may lead to compounded errors, as inaccuracies generated in earlier timesteps accumulate over time, resulting in inconsistent or divergent sequences that deviate from the true distribution of the target sequence. Thirdly, autoregressive models rely on a left-to-right causal attention mechanism, limiting the expressive power of hidden representations, as each item encodes information solely from preceding items(Sun et al., 2019). This constraint impedes optimal representation learning, resulting in suboptimal performance.